OpenStudy (julierosehead):

Simplify the expression 5(x-z) + 4(y-x) I came up with 9(z+y) - (x2) X to the second power by the way I don't know if I'm correct or not

OpenStudy (mathmale):

Would you mind typing in the procedure you used to obtain your 9(z+y)-x^2 result, please? Then I could give you better feedback. Note: "the square of x" is written as x^2, not as x2.

OpenStudy (mathmale):

5(x-z)=?

OpenStudy (julierosehead):

Ok thanks lol I didn't know how to type it

OpenStudy (julierosehead):

This is what I did, like I said I don't really know how to do it.. 5(x-y) + 4(y-x) = 9 • (x-z) + (y-x) = 9(z+y - x^2)

OpenStudy (julierosehead):

Please help me it's one of the last questions I have in this exam